In the energy transformation in multiple collisions experiment, if the initial release height is 175 cm and the first rebound height is 155 cm what would you expect to measure as the 5th rebound height?
Answer #1

let

H = 175 cm
h1 = 155 cm

coefficient of restitution, e = h1/H

e = 155/175

= 0.8857
so,
h2/h1 = e
h3/h2 = e
h4/h3 = e
h5/h4 = e

now use,
(h5/h4)*(h4/h3)*(h3/h2)**(h2/h1)*(h1/H) = e^5

h5/H = e^5

h5 = e^5*H

= 0.8857^5*175

= 95.4 cm <<<<<<<<<<<----------------Answer